OpenStudy (anonymous):

NO GRAPHING CALC, HELP PLZ Solve the equation by graphing: 5x^3 − 7x^3 − 20x = 0

OpenStudy (anonymous):

Hey did you really mean to put x cubed on two of them?

OpenStudy (anonymous):

yes i did.

OpenStudy (anonymous):

oh so then we can simplify to \[-2x^3-20x=0\]

OpenStudy (anonymous):

i do not have graphing calc

OpenStudy (anonymous):

=>\[ -2x(x^2+10) \]

OpenStudy (anonymous):

it has x-intercepts at x= 0, \[-\sqrt10, \sqrt 10\]

OpenStudy (anonymous):

? Answer choices -1.42, 0, 2.82 0 -28.9, 0, 2.82 No solutions

OpenStudy (anonymous):

what does this have to do with graphing? What is the exact question? Is there two equations?

OpenStudy (anonymous):

this....Solve the equation by graphing: 5x^3 − 7x^3 − 20x = 0

OpenStudy (anonymous):

ok then did they want you to plug it in a calculator?

OpenStudy (anonymous):

because if they did then from what you are saying it would be positive or negative 3.16. I think though that it is more likely that the middle should be x^2

OpenStudy (anonymous):

thts the question and those r the answer choiices..

OpenStudy (anonymous):

x= -3.16, 0, 3.16

OpenStudy (anonymous):

but no. mayb your right

OpenStudy (anonymous):